# Large Language Models meet Collaborative Filtering: An Efficient All-round LLM-based Recommender System

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 5.1. Experimental Setup
Datasets. For comprehensive evaluations, we used four datasets from Amazon datasets (McAuley et al., 2015b; He and McAuley, 2016), i.e., Movies and TV, Video Games, Beauty, and Toys, which consist of comprehensive textual information including ”title” and ”description.” Note that we deliberately selected datasets with varying statistics in terms of number of users and items to conduct an extensive analysis of the models. The statistics for each dataset after preprocessing are presented in Table 10 in the Appendix B. For more details regarding the datasets and how they are preprocessed, please refer to Appendix B.
Baselines. We compare A-LLMRec with the following baselines that can be categorized into three types: collaborative filtering recommender systems (NCF (He et al., 2017), NextItNet (Yuan et al., 2019), GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018)), modality-aware recommender systems (MoRec (Yuan et al., 2023), CTRL (Li et al., 2023a), and RECFORMER (Li et al., 2023b)), and LLM-based recommender systems (LLM-Only, TALLRec (Bao et al., 2023) and MLP-LLM). For more detail regarding the baselines, please refer to Appendix C
@@
Implementation Details. Although A-LLMRec is model-agnostic, in this work, we adopt OPT-6.7B (Zhang et al., 2022) as the backbone LLM and SASRec (Kang and McAuley, 2018) as the pre-trained CF-RecSys. For fair comparisons, we also used OPT-6.7B as the backbone LLM for other LLM-based models (i.e., LLM-Only, TALLRec (Bao et al., 2023) and MLP-LLM). Moreover, we use SASRec as the CF-RecSys in other modality-aware models (i.e., MoRec (Yuan et al., 2023) and CTRL (Li et al., 2023a)), and fix the dimension of item and model embeddings to 50 for all the methods and datasets. For RECFORMER (Li et al., 2023b), we follow the paper and employ Longformer (Beltagy et al., 2020) as the backbone network. We set the batch size to 128 for all collaborative filtering-based and modality-aware models. Moreover, the batch size is set to 32 for Stage-1 of A-LLMRec, and 4 for MLP-LLM, TALLRec, and Stage-2 of A-LLMRec. We trained Stage-1 of A-LLMRec for 10 epochs, and Stage-2 of A-LLMRec for 5 epochs, and TALLRec is trained for a maximum of 5 epochs. We use the Adam optimizer to train the models in all datasets. For hyperparameters, we tune the model in certain ranges as follows: learning rate $\eta_{1},\eta_{2}$ in $\left\{0.01,0.001,0.0005,0.0001\right\}$ for the training stage each, coefficient $\alpha,\beta$ in $\left\{0.1,0.2,0.5,0.75,1.0\right\}$ for each. We report the best-performing hyper-parameters for each dataset in Table  11 of Appendix D. We use four NVIDIA GeForce A6000 48GB for the Movies and TV dataset to train LLM-based models, and one NVIDIA GeForce A6000 48GB for other datasets including LLM-based and other models.
@@ 5.3.2. Effect of the Alignment method in Stage-2
Recall that a user representation and item embeddings are injected to the LLM prompt as shown in Figure 3. In this section, we verify the benefit of injecting them into the prompt (rows (2-4) in Table  7). We have the following observations: Across all datasets, 1) the absences of either the user representation (row (2)) or the joint embedding (row (3)) from the prompt led to a reduction in performance. Notably, the exclusion of the joint embedding results in a more substantial decrease, underscoring its significant role in transferring collaborative knowledge. Moreover, as joint embeddings also capture the textual information about items, their exclusion is particularly detrimental. 2) When we replace the joint embedding with a randomly initialized embedding (row (4)), which means A-LLMRec is trained with item embeddings without collaborative knowledge, we observe performance degradation across all datasets. This indicates the importance of leveraging the collaborative knowledge for recommendation.
@@ 5.4.4. Beyond Recommendation: Language Generation Task (Favorite genre prediction)
To validate whether A-LLMRec can generate natural language outputs based on the understanding of users and items through the aligned collaborative knowledge from CF-RecSys, we conduct a favorite genre prediction task (Figure 4). That is, given the same prompt format, we ask the LLM-based models (i.e., A-LLMRec and LLM-Only) using the same backbone LLM, which is OPT-6.7B, to predict the movie genres that a given user would enjoy watching. The only difference in the prompt is that while LLM-only is only given titles of movies watched by the user in the past, A-LLMRec is given the user representation and item embeddings along with the movie titles. In Figure 4, we observe that A-LLMRec indeed generates proper answers, while LLM-Only fails to do so. We attribute this to the fact that the item embeddings of the CF-RecSys are well aligned with the token space of the LLM, which enables the LLM to understand and utilize collaborative knowledge. Note that although we also experimented with TALLRec, we were not able to obtain valid outputs. We conjecture that since the LLM in TALLRec is fine-tuned via an instruction-tuning process that makes the model provide responses as part of the recommendation task, generating valid natural language outputs has become a non-trivial task. Please refer to Appendix E for the results of TALLRec.
@@ 6. Conclusion
Appendix A Ethics Statement
To the best of our knowledge, this paper aligns with the KDD Code of Ethics without any ethical concerns. The datasets and codes employed in our research are publicly available.
Appendix B Datasets
Datasets #Users #Items #Interactions. Avg. Len
Movies and TV 297,498 59,944 3,409,147 11.46
Video Games 64,073 33,614 598,509 8.88
Beauty 9,930 6,141 63,953 6.44
Toys 30,831 61,081 282,213 9.15
Table 10 shows the statistics for each dataset after preprocessing, and we describe details regarding data preprocessing as follows:
Appendix C Baselines
@@
    • NCF (He et al., 2017) combines neural networks (MLP) to capture the collaborative information. Note that NCF is a two-tower model, which is comprised of separate components for the user and item embedding matrix.
    • NextItNet (Yuan et al., 2019) proposes a temporal convolutional network that utilizes 1D-dilated convolutional layers to capture the long-term dependencies inherent in interaction sequence, and by utilizing residual connections, NextItNet enhances optimization efficiency.
@@
    • CTRL (Li et al., 2023a) employs a two-stage learning process: the first stage involves contrastive learning on textual information of items to initialize the backbone model, and the second stage, fine-tune the model on recommendation tasks. We use SASRec as the backbone model of CTRL.
@@
    • TALLRec (Bao et al., 2023) is our main baseline, which learns the recommendation task based on prompts consisting solely of text and fine-tunes the LLMs using the LoRA. Their approach involves providing user interaction history and one target item and determining whether a user will prefer this target item. This simpler task necessitates only a brief prompt for the LLMs. The authors can set a maximum input length of 512 and a batch size of 128. In contrast, our recommendation task requires a more extensive prompt, more than twice as long, as it must include 20 candidate items. Consequently, we have disabled the truncation mode for the LLM and removed the maximum length constraint, even though this adjustment results in a smaller batch size, the same as  A-LLMRec, for training TALLRec. We use the prompt shown in Figure 5.
@@
Figure 5. An example prompt designed for the Amazon Movies dataset used by LLM-based models, i.e., TALLRec and LLM-Only models. Note that the prompt for MLP-LLM and A-LLMRec is shown in Figure 3, which only differs in providing user representation and item embeddings.
Table 11. Hyperparameter specifications of  A-LLMRec
Learning rate Learning rate embedding dim embedding dim alpha beta
stage 1 stage 2 (CF-RecSys) $d$ ($f_{I}^{enc},f_{T}^{enc}$) $d^{\prime}$
Movies and TV 0.0001 0.0001 50 128 0.5 0.5
Video Games 0.0001 0.0001 50 128 0.5 0.5
Beauty 0.0001 0.0001 50 128 0.5 0.2
Toys 0.0001 0.0001 50 128 0.5 0.2
Figure 6. A-LLMRec, LLM-Only, and TALLRec on the favorite genre prediction task (Movies and TV dataset used).
Table 12. Stage-1 hyperparamter anaylsis (Hit@1).
$\alpha$ $\beta$ Beauty
 A-LLMRec 0.5 0.2 0.5809
1 1 0.5656
0.5 0.5 0.5722
0.2 0.2 0.5688
TALLRec - 0.5542
@@
Methods Source code
SASRec https://github.com/pmixer/SASRec.pytorch
NextItNet https://github.com/syiswell/NextItNet-Pytorch
GRU4Rec https://github.com/hungpthanh/GRU4REC-pytorch
RECFORMER https://github.com/AaronHeee/RecFormer
TALLRec https://github.com/SAI990323/TALLRec
A-LLMRec https://github.com/ghdtjr/A-LLMRec
Appendix D Hyperparamter Analysis
In Table 12 we conduct an analysis regarding the hyperparameter sensitivity on the Beauty dataset to verify the robustness of A-LLMRec over hyperparameters, especially on $\alpha$ and $\beta$ in Equation 6. In Table  12, although the performance of A-LLMRec is the best when $\alpha=0.5$ and $\beta=0.2$A-LLMRec still outperforms TALLRec without carefully setting the hyperparameters. This implies the superiority of A-LLMRec.
Appendix E Language Generation Task
In Figure 6, we present additional favorite genre prediction task results for experiment in shown in Section 5.4.4. As mentioned in Section 5.4.4, TALLRec could not generate valid natural language outputs due to the fine-tuning via instruction tuning process, which makes the LLM of TALLRec being able to answer only with some particular prompts used in instruction tuning process. Moreover, the additional results also indicate that A-LLMRec can generate the favorite genres for the users based on the deep understanding of the aligned user representation and item embeddings while LLM-only fails to do so.
Appendix F Reproducibility
For implementing the baseline, we followed the official codes published by authors as detailed in Table 13. We conduct experiments within the same environment. Refer to our source code and instructions to run code for reproducing the results reported in the experiments.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/XXXXXX.XXXXXX
@@ 4.2.2. Prompt Design for Integrating Collaborative Knowledge
Datasets #Users #Items #Interactions. Avg. Len
Movies and TV 297,498 59,944 3,409,147 11.46
Video Games 64,073 33,614 598,509 8.88
Beauty 9,930 6,141 63,953 6.44
Toys 30,831 61,081 282,213 9.15
@@ 5.1. Experimental Setup
Datasets. For comprehensive evaluations, we used four datasets from Amazon datasets (McAuley et al., 2015b; He and McAuley, 2016), i.e., Movies and TV, Video Games, Beauty, and Toys, which consist of comprehensive textual information including ”title” and ”description.” Note that we deliberately selected datasets with varying statistics in terms of number of users and items to conduct an extensive analysis of the models. The statistics for each dataset after preprocessing are presented in Table 2 and we describe details regarding data preprocessing as follows:
Baselines. We compare A-LLMRec with the following baselines that can be categorized into three types: collaborative filtering recommender systems (NCF (He et al., 2017), NextItNet (Yuan et al., 2019), GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018)), modality-aware recommender systems (MoRec (Yuan et al., 2023), CTRL (Li et al., 2023a), and RECFORMER (Li et al., 2023b)), and LLM-based recommender systems (LLM-Only, TALLRec (Bao et al., 2023) and MLP-LLM). For more detail regarding the baselines, please refer to Appendix A
Table 3. Hyperparameter specifications of  A-LLMRec
Learning rate Learning rate embedding dim embedding dim alpha beta
stage 1 stage 2 (CF-RecSys) $d$ ($f_{I}^{enc},f_{T}^{enc}$) $d^{\prime}$
Movies and TV 0.0001 0.0001 50 128 0.5 0.5
Video Games 0.0001 0.0001 50 128 0.5 0.5
Beauty 0.0001 0.0001 50 128 0.5 0.2
Toys 0.0001 0.0001 50 128 0.5 0.2
@@
Implementation Details. Although A-LLMRec is model-agnostic, in this work, we adopt OPT-6.7B (Zhang et al., 2022) as the backbone LLM and SASRec (Kang and McAuley, 2018) as the pre-trained CF-RecSys. For fair comparisons, we also used OPT-6.7B as the backbone LLM for other LLM-based models (i.e., LLM-Only, TALLRec (Bao et al., 2023) and MLP-LLM). Moreover, we use SASRec as the CF-RecSys in other modality-aware models (i.e., MoRec (Yuan et al., 2023) and CTRL (Li et al., 2023a)), and fix the dimension of item and model embeddings to 50 for all the methods and datasets. For RECFORMER (Li et al., 2023b), we follow the paper and employ Longformer (Beltagy et al., 2020) as the backbone network. We set the batch size to 128 for all collaborative filtering-based and modality-aware models. Moreover, the batch size is set to 32 for Stage-1 of A-LLMRec, and 4 for MLP-LLM, TALLRec, and Stage-2 of A-LLMRec. We trained Stage-1 of A-LLMRec for 10 epochs, and Stage-2 of A-LLMRec for 5 epochs, and TALLRec is trained for a maximum of 5 epochs. We use the Adam optimizer to train the models in all datasets. For hyperparameters, we tune the model in certain ranges as follows: learning rate $\eta_{1},\eta_{2}$ in $\left\{0.01,0.001,0.0005,0.0001\right\}$ for the training stage each, coefficient $\alpha,\beta$ in $\left\{0.1,0.2,0.5,0.75,1.0\right\}$ for each, we report the best-performing hyper-parameters for each dataset in Table  3. We use four NVIDIA GeForce A6000 48GB for the Movies and TV dataset to train LLM-based models, and one NVIDIA GeForce A6000 48GB for other datasets including LLM-based and other models.
@@ 5.3.2. Effect of the Alignment method in Stage-2
Recall that a user representation and item embeddings are injected to the LLM prompt as shown in Figure 3. In this section, we verify the benefit of injecting them into the prompt (rows (2-4) in Table  9). We have the following observations: Across all datasets, 1) the absences of either the user representation (row (2)) or the joint embedding (row (3)) from the prompt led to a reduction in performance. Notably, the exclusion of the joint embedding results in a more substantial decrease, underscoring its significant role in transferring collaborative knowledge. Moreover, as joint embeddings also capture the textual information about items, their exclusion is particularly detrimental. 2) When we replace the joint embedding with a randomly initialized embedding (row (4)), which means A-LLMRec is trained with item embeddings without collaborative knowledge, we observe performance degradation across all datasets. This indicates the importance of leveraging the collaborative knowledge for recommendation.
@@ 5.4.4. Beyond Recommendation: Language Generation Task (Favorite genre prediction)
To validate whether A-LLMRec can generate natural language outputs based on the understanding of users and items through the aligned collaborative knowledge from CF-RecSys, we conduct a favorite genre prediction task (Figure 4). That is, given the same prompt format, we ask the LLM-based models (i.e., A-LLMRec and LLM-Only) using the same backbone LLM, which is OPT-6.7B, to predict the movie genres that a given user would enjoy watching. The only difference in the prompt is that while LLM-only is only given titles of movies watched by the user in the past, A-LLMRec is given the user representation and item embeddings along with the movie titles. In Figure 4, we observe that A-LLMRec indeed generates proper answers, while LLM-Only fails to do so. We attribute this to the fact that the item embeddings of the CF-RecSys are well aligned with the token space of the LLM, which enables the LLM to understand and utilize collaborative knowledge. Note that although we also experimented with TALLRec, we were not able to obtain valid outputs. We conjecture that since the LLM in TALLRec is fine-tuned via an instruction-tuning process that makes the model provide responses as part of the recommendation task, generating valid natural language outputs has become a non-trivial task. Please refer to Appendix B for the results of TALLRec.
@@ 6. Conclusion
Ethics Statement To the best of our knowledge, this paper aligns with the KDD Code of Ethics without any ethical concerns. The datasets and codes employed in our research are publicly available.
Acknowledgements.
This work was supported by NAVER Corporation, the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00335098), and National Research Foundation of Korea(NRF) funded by Ministry of Science and ICT (NRF-2022M3J6A1063021).
Appendix A Baselines
@@
    • NCF (He et al., 2017) combines neural networks (MLP) to capture the collaborative information. Note that NCF is a two-tower model comprised of separate components for the user and item embedding matrix.
    • NextItNet (Yuan et al., 2019) proposes a temporal convolutional network that utilizes 1D-dilated convolutional layers and residual connections to capture the long-term dependencies inherent in interaction sequence.
@@
    • CTRL (Li et al., 2023a) employs a two-stage learning process: the first stage involves contrastive learning on textual information of items to initialize the backbone model, and the second stage, fine-tunes the model on recommendation tasks. We use SASRec as the backbone model of CTRL.
@@
    • TALLRec (Bao et al., 2023) is our main baseline, which learns the recommendation task based on prompts consisting solely of text and fine-tunes the LLMs using the LoRA. Their approach involves providing user interaction history and one target item and determining whether a user will prefer this target item. This simpler task necessitates only a brief prompt for the LLMs. In contrast, our recommendation task requires a more extensive prompt. Even though this adjustment results in a smaller batch size, the same as  A-LLMRec, for training TALLRec. We use the prompt shown in Figure 6.
@@
Figure 6. An example prompt designed for the Amazon Movies dataset used by LLM-based models, i.e., TALLRec and LLM-Only models.
@@
Appendix B Language Generation Task
In Figure 5, we present additional favorite genre prediction task results for experiment in shown in Section 5.4.4. As mentioned in Section 5.4.4, TALLRec could not generate valid natural language outputs due to the fine-tuning via instruction tuning process, which makes the LLM of TALLRec being able to answer only with some particular prompts used in instruction tuning process. The additional results indicate that A-LLMRec can generate the favorite genres for the users based on the understanding of the aligned user representation and item embeddings while LLM-only fails to do so.
Appendix C Reproducibility
For implementing the baseline, we followed the official codes published by authors as detailed in Table 12. Refer to our source code and instructions to run code for reproducing the results reported in the experiments.